OpenStudy (anonymous):

The postman measures a large envelope to determine the correct postage. He knows the area of the envelope, A, can be determined with the equation A = lw, where l is the length and w is the width. Which equation can he use to solve for the length? l = Aw l = A ÷ w l = w + A l = A – w

OpenStudy (anonymous):

What's the problem? You need to have l on one side and the rest on the other. What could you do to get rid of that w on the left side?

OpenStudy (anonymous):

not so sure

OpenStudy (anonymous):

What's opposite to multiplication?

OpenStudy (anonymous):

divid but im not surte what to divid

OpenStudy (anonymous):

Okay, let's change that equation into something easier. Let's say you have 2l = 10 and you want to calculate l. What would you do? What is l?

OpenStudy (anonymous):

try to get l

OpenStudy (anonymous):

Yeah, and how? If 2 * l = 10, what's the value of l?

OpenStudy (anonymous):

is it 4

OpenStudy (anonymous):

No, because 2*4 = 8. The answer is 5. But anyway, how did you try to get this answer? Did you just guess? You know, when you have 2 boxes of apples and you know there is the same amount of apples in each box, and you also know there are 10 apples, would you know how many apples are in one box?

OpenStudy (anonymous):

no i got 4 because i multpled wrong i got it wrong

OpenStudy (anonymous):

but 2 time 4 is 8 i mad a easy mistakr

OpenStudy (anonymous):

Yeah, I noticed that. I would recommend you to really brush up on those basic skills. It would be very hard, trying to understand any more advanced math without such knowledge. And okay, I'll tell you what to do. You know to divide both sides by w. Then you get: A / w = l * w/w. And w/w = 1. So you can you get what you wanted: A/w = l.

OpenStudy (anonymous):

im good with the other math just the letter parts i dont know

OpenStudy (anonymous):

Unfortunately, the more you learn math, the more letters it uses.

OpenStudy (anonymous):

is the answer I=AW

OpenStudy (anonymous):

\[l = A \div w\]

OpenStudy (anonymous):

OHHH THANKS

OpenStudy (anonymous):

I UNDERSTAND THANK YOU SO MUCH